### v3

# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

Abstract
Supply chain risk assessment (SCRA) has witnessed a profound evolution through the integration of artificial intelligence (AI) and machine learning (ML) techniques, revolutionizing predictive capabilities and risk mitigation strategies. The significance of this evolution stems from the critical role of robust risk management strategies in ensuring operational resilience and continuity within modern supply chains. Previous reviews have outlined established methodologies but have overlooked emerging AI/ML techniques, leaving a notable research gap in understanding their practical implications within SCRA. This paper conducts a systematic literature review combined with a comprehensive bibliometric analysis. We meticulously examined 1,439 papers and derived key insights from 51 articles published between 2015 and 2024. The review fills this research gap by addressing pivotal research questions and exploring existing AI/ML techniques, methodologies, findings, and future trajectories, thereby providing a more encompassing view of the evolving landscape of SCRA. Our study unveils the transformative impact of AI/ML models, such as Random Forest, XGBoost, and hybrids, in substantially enhancing precision within SCRA. It underscores adaptable post-COVID strategies, advocating for resilient contingency plans and aligning with evolving risk landscapes. Significantly, this review surpasses previous examinations by accentuating emerging AI/ML techniques and their practical implications within SCRA. Furthermore, it highlights the contributions through a comprehensive bibliometric analysis, revealing publication trends, influential authors, and highly cited articles. This analysis provides an encompassing view of the field’s growth, offering insights into its evolution. This comprehensive review provides a roadmap for practitioners and researchers, offering insights into fortifying SCRM strategies through AI integration, ultimately contributing to a deeper understanding of evolving trends and applications in this dynamic field.
keywords:
Artificial Intelligence (AI) , Supply chain risk assessment , Bibliometric analysis , Content analysis , Systematic literature review
[…]
[inst3]organization=Department of Computer Science,addressline=American International University-Bangladesh (AIUB), city=Dhaka, postcode=1229, country=Bangladesh
1 Introduction
Supply chain management (SCM) has long grappled with various challenges from various sources. Past outbreaks of infectious diseases, geological upheavals like earthquakes, and other natural catastrophes have put supply chains (SCs) at risk, albeit on a limited scale (Govindan et al.,, 2020). These incidents have illustrated how interrelationships within SCs have formed intricate risk contagion networks, making them susceptible to contagion effects (Agca et al.,, 2022). The cascade effect, characterized by risk spillover from one enterprise to another, exacerbates the challenges, culminating in a chain reaction of SC Resilience (SCR) disasters (Roukny et al.,, 2018).
Across the globe, companies confront formidable challenges across all stages of their SCs. Suppliers failing to meet delivery obligations, unpredictable shifts in customer demands, and episodes of panic buying are just a few examples of the hurdles companies face (Ivanov,, 2020). In pursuing market leadership, enterprises must also contend with the complexity introduced by the management and launch of innovation projects, further complicating SCM (Kwak et al.,, 2018). Furthermore, digital transformation has ushered in a wave of technological advancements in SC operations (Kwak et al.,, 2018).
However, the most significant disruption in recent times emerged during the COVID-19 pandemic, impacting global SCs profoundly. The disruptions encompassed not only the movement of people but also the flow of raw materials and finished goods, along with extensive disruptions in factory and SC operations (Sheng et al.,, 2021). These disruptions ushered in unprecedented challenges for SC professionals as they grappled with an entirely new reality (Araz et al.,, 2020; Craighead et al.,, 2020). The pandemic represents a distinctive form of SC disruption, distinct from natural or man-made disasters (e.g., the Japanese earthquake and 9/11 attacks) and disruptions driven by evolving technologies and changing customer attitudes (Zhang et al.,, 2020).
While existing guidance exists for predicting, managing, and responding to these disruptions, the challenges brought about by COVID-19 have underscored the paramount importance of risk management, unlike any previous disruption (Bode et al.,, 2011; Craighead et al.,, 2020). Consequently, the concept of SCR in the context of COVID-19 has emerged as a critical area necessitating further exploration and development.
SCs are not the sole victims in this precarious environment, as the ripple effect extends to upstream and downstream enterprises. Practical risk management programs are imperative for enterprises to avert SCR and enhance SC risk management (SCRM). Moreover, incorporating advanced AI technologies, such as ML, can provide predictive capabilities for SCRM. ML algorithms have showcased their ability to identify abnormal risk factors and derive predictive insights from historical data (Guo et al.,, 2021; Mohanty et al.,, 2021). By harnessing ML, enterprises can detect risk factors and anticipate market demands and potential risk scenarios (Punia et al.,, 2020; Wu et al., 2022a, ). ML’s proficiency in handling non-linear relationships further bolsters its superiority over traditional linear models. Additionally, its aptitude for processing unstructured data, a challenge where traditional models falter, positions ML as a formidable tool for addressing time, cost, and resource constraints within SCs.
However, despite its immense potential, SC researchers have historically exhibited limited familiarity with ML compared to other SCM aspects like mathematical programming and stochastic optimization (Liu et al.,, 2019; Shahed et al.,, 2021). Furthermore, the classification of ML algorithms remains unclear (Janiesch et al.,, 2021; Xu and Jackson,, 2019). Bridging this knowledge gap and exploring the value of ML for SCRM through interdisciplinary integration is a critical research need. Remarkably, there has been no prior effort to scrutinize the SCRM literature within an ML environment, in stark contrast to previous reviews primarily centered on risk definition, classification, and management strategies.
Companies should augment their analytical capabilities by harnessing organizational knowledge to enhance SCR, thereby elevating their information capabilities (Wong et al.,, 2020). As underscored by previous studies, the role of AI-enabled technologies extends to promoting innovations for enhanced SC performance (SCP) (Baryannis et al., 2019b, ; Nayal et al.,, 2021). The adaptation capabilities and information processing prowess offered by AI techniques hold the potential to enhance SCP (Belhadi et al., 2021a, ). Notably, AI has found application across diverse sectors, improving flexibility and communication and reducing undue fluctuations for successful project execution (Lalmi et al.,, 2021). To navigate the impact of risks and disruptions, (Katsaliaki et al.,, 2022) advocates integrating three critical facets: long-term partnerships, IT applications for business enhancement, and government policies that facilitate adaptability.
The contemporary SC landscape demands the integration of AI, specifically ML, to invigorate SCRA and SCRM. While traditional methods possess merit, their limitations can be effectively addressed by AI, which offers predictive capabilities, nonlinear relationship analysis, and unstructured data processing. As SCs continue to evolve, AI-based risk assessment is imperative for fostering resilience and sustaining the efficiency and effectiveness of SC operations. Existing reviews of SCRA using ML often fail to offer a comprehensive view of evolving ML techniques, neglecting emerging trends and overlooking diverse applications and non-journal sources. They also tend to overlook the role of ML in the response phases of SCRM and do not sufficiently address data-related challenges.
In light of the limitations associated with traditional SCRA methods, there is a growing interest in harnessing the potential of AI techniques to enhance risk assessment practices. This paper presents a systematic literature review (SLR) focused on the application of AI in SCRA. The existing literature lacks bibliometric analysis incorporating recent and adequate articles on the AI-SCRA topic. Another limitation is that most literature only contributed to bibliometric or content analysis. We tried to close this gap by including both analyses in our study, leveraging the systematic review methodology. The main objective of this review is to analyze the existing literature critically, identify research gaps, and provide insights into the use of AI techniques, such as ML, Deep Learning (DL), and natural language processing (NLP), to improve the accuracy and effectiveness of SCRA. Our paper has the following contributions-
  1. 1.
    Our study provides a comprehensive review of AI and ML applications in SCRA, supported by bibliometric and co-citation analyses to capture recent advancements and trends in the field.
  2. 2.
    We conduct content analysis to explore thematic trends, detailing the advantages, disadvantages, and specific application scenarios of various AI/ML models used in SCRA, offering insights into their suitability for different SC contexts.
  3. 3.
    Our proposed AI-based framework guides practitioners in assessing SC risks, from data collection to model selection and interpretation, facilitating efficient risk management in dynamic SCs.
The rest of the article is organized as follows: the “Literature Review” examines the current body of knowledge. The “Methodology” section details the study’s design, including search strategies and selection criteria. The “Bibliometric Analysis” provides insights from bibliometric tools, with visualizations and observations on AI in SCRA. A structured framework for SCRA is given in “Framework for SCRA using AI section. In the “Content Analysis of AI Techniques in SCRA,” the review discusses various AI techniques, highlighting their benefits, limitations, and real-world applications. Practical insights for industry are provided in the “Practical and Managerial Implications” section, offering actionable guidance for practitioners. The review also outlines key “Challenges and Limitations” in this area, and suggests “Future Research Directions” for ongoing exploration. The paper concludes with a summary of findings and recommendations in the “Conclusion.”
2 Literature Review
Deiva Ganesh and Kalpana, (2022) conducted a comprehensive and descriptive literature study to identify AI and ML approaches in SCRM stages. This study examined SCRM research publications from three scientific databases from 2010 to 2021. They proposed a data analytics, simulation, and optimization framework that might produce a comprehensive SC risk identification, assessment, mitigation, and monitoring strategy. Hence, leveraging AI, Blockchain, and the Industrial Internet of Things (IIoT) to create smarter SCs can transform how firms handle uncertainty. This analysis of Deiva Ganesh and Kalpana, (2022) does not evaluate blockchain, big data, and IIoT-related SC articles, which may limit information. Due to limited research, only recent English-language publications on SCRM and AI were included. Nimmy et al., (2022) did a thorough literature evaluation on operational risk assessment methods and whether AI explains them. The report suggested that risk managers use auditable SC operational risk management methodologies to understand why they should take a risk management action rather than just what to do. While they employed the SLR method, they could only examine some AI solutions for SCRM. They suggested hybridizing SC operational risk management and explainable AI (XAI) to obtain XAI-like characteristics in future research. Li et al., (2023) analyzed COVID-19 SCR research history, present, and future. In particular, supervised ML classifies 1717 SCR papers into 11 subject categories. Each cluster was then studied in the context of COVID-19, indicating three related skills (interconnectedness, transformability, and sharing) on which enterprises could work to develop a more resilient SC post-COVID. Their data only came from Scopus’ core collection, which might affect the results; thus, they suggested adding WoS and EBSCO to the evaluation (Xu et al.,, 2020). Given the fast expansion of SCR research, they only picked English-language articles, which may have excluded valuable knowledge. They recommended network analysis to determine cluster linkages and SCR literary themes.
Naz et al., (2022) examined the role of AI in building a robust and sustainable SC and offered optimal risk mitigation options. For review, 162 SCOPUS research publications were selected. Based on the nominated articles, Structural Topic Modeling produced various AI-related theme topics in SCR. AI research trends in SCR were examined using R-package bibliometric analysis. They solely studied journal articles, not conference papers, field reports, corporate reports, book chapters, etc. Ni et al., (2020) analyzed articles from 1998/01/01 to 2018/12/31 in five major databases to highlight the newest research trends in the field. ML applications in SCM were still developing due to a lack of high-yielding authors and poor publication rates. 10 ML algorithms were extensively utilized in SCM, but their utilization was unequal among the SCM tasks most often reported. This paper has limitations in reviewing only five popular databases to limit articles for evaluation that might have filtered some related articles. Second, only widely used ML methods in SCM were counted in this review, and other ML techniques brought to SCM may be helpful later on. Low-frequency ML algorithms should be analyzed for further research in this area. Finally, their article contained 32 well-recognized ML methods; however, some newly generated ML algorithms may be used in SCM after 2015. Baryannis et al., 2019b examined SC risk definitions and uncertainty. Then, a mapping analysis categorizes available literature by AI approaches and SCRM tasks. Most of the examined works focus on building and assessing a mathematical model that accounts for various uncertainties and hazards but less on establishing and analyzing the applicability of the suggested models. They found that only 9 of 276 research (3%) use comprehensive techniques that cover all three SCRM stages (identification, assessment, and response). Yang et al., (2023) thoroughly examined the advancement of ML algorithms in SCRM by gathering 67 publications from 9 authoritative databases in the first half of 2021. They analyzed only English language journal articles from 9 relevant academic databases, excluding conference papers, textbooks, and unpublished articles and notes. This analysis relied on keyword searches, which may have missed some work. Schroeder and Lodemann, (2021) used a comprehensive and multi-vocal literature study to get a complete picture of our subject field. The SLR identified 533 papers in this area and examined 23. The comprehensive literature study found that just a few examples of ML in SCRM have been reported in depth in the scientific literature, and those examples focus on manufacturing, transport, and the complete SC. Shi et al., (2022) examined 76 works on credit risk utilizing statistical, ML, and DL methods over the last eight years. They offered a unique classification approach and performance rating for ML-driven credit risk algorithms utilizing public datasets. Data imbalance, dataset inconsistency, model transparency, and DL model underuse are discussed. Their review found that most DL models outperform standard ML and statistical algorithms in credit risk prediction, and ensemble techniques outperform single models.
Previous studies in SCRM have shown limitations in evaluating certain technologies, such as blockchain, big data, and the IIoT, potentially limiting the scope of analysis. Our review addresses this gap by incorporating an analysis of these technologies within the context of SCRM, providing a more comprehensive assessment. There is a recognized need in the literature to integrate SC operational risk management with XAI for future research. Our review contributes to closing this gap by exploring the integration of these approaches and discussing their implications for SCRM. Some studies have highlighted the importance of utilizing multiple databases to evaluate SCRM literature comprehensively. Our review extends this by considering insights from the most reliable and prominent two databases, enhancing the breadth of our analysis. Previous research has found that only a small percentage of studies cover all three stages of SCRM comprehensively. Our review contributes to closing this gap by providing a holistic assessment of SCRM stages, including identification, assessment, and response strategies, leveraging insights from diverse studies.
3 Methodology
3.1 Study Design
The research utilized a systematic approach to conduct a comprehensive literature review and a bibliometric analysis. This methodology aimed to scrutinize the intersection of AI/ML techniques and SCRA, providing insights into research trends, methodologies, and emerging themes. We formulated the following research questions to guide our SLR:
  • RQ1:
    What is the current state of research on applying AI/ML techniques in SCRM?
  • RQ2:
    Which AI/ML techniques are commonly employed in SCRA, prediction, and mitigation?
  • RQ3:
    What are the key findings and trends identified in the literature regarding using AI/ML in SCRM?
  • RQ4:
    What are the research gaps and future directions in this field?
The SLR followed a robust methodology, adhering to the guidelines proposed by Denyer and Tranfield, (2009). This approach aimed to identify and analyze research articles focusing on integrating AI/ML techniques within SCRA.
3.2 Search Strategy
We conducted a comprehensive search using two major academic databases: Google Scholar and Web of Science (WoS). We used the “Publish or Perish 8” (Harzing,, 2016) software for the Google Scholar search to filter articles published between 2015 and 2024. The initial search yielded 1000 articles. In the WoS search, we employed the following search strings to retrieve relevant articles:
(ALL=( “supply chain”) AND (“risk*” OR “credit risk*” OR “risk assessment*” OR “risk prediction” OR “disruption*” OR “disease outbreak*” OR “post COVID resilienc*”) AND (“artificial intelligence” OR “machine learning” OR “neural network*” OR “deep learning” OR “reinforcement learning” OR “SVM” OR “support vector machine” OR “boosting” OR “ensemble” OR “bayesian network model” OR “random forest” OR “LSTM” OR “long short-term memory”))) NOT (DT==(“PROCEEDINGS PAPER” OR “BOOK CHAPTER” OR “EDITORIAL MATERIAL” OR “LETTER”))
Table 1: Inclusion and exclusion criteria and justifications
Criterion Inclusion/Exclusion Justification
Date range 2015–2024 Focuses on recent studies, ensuring relevance to current advancements in AI for SCRA.
Language English only Ensures consistency in analysis and accessibility for an English-speaking academic audience.
Duplicate removal Exclusion Avoids redundancy and ensures each study contributes uniquely to the analysis.
Publisher reliability Exclusion Maintains academic quality by including only studies from reputable, peer-reviewed journals.
Publication type Exclusion of Q3/Q4, proceedings, book chapters, etc. Focuses on high-impact research, prioritizing quality over quantity.
Checking for relevance (title, abstract, and conclusion) Exclusion Ensures alignment with the study’s objectives based on content relevant to SC risk and AI.
3.3 Article Selection Process
To ensure the selection of high-quality and relevant studies, we followed a systematic, multi-stage screening process guided by the PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) framework. The selection process was designed to minimize potential bias and ensure transparency at each stage. The stages are detailed as follows:
  1. 1.
    Removal of Duplicates: To avoid redundant analysis, we first identified and removed duplicate articles between the two databases, excluding 360 duplicate entries. This step is standard in systematic reviews to ensure each study contributes uniquely to the analysis.
  2. 2.
    Filtering by Publisher Reliability: We excluded 398 articles published by sources identified as less reliable, focusing only on those published in reputable, peer-reviewed journals. This criterion helps maintain high academic integrity and ensures that the final selection is based on studies that meet established research quality standards.
  3. 3.
    Non-English Exclusion: Since this review focuses on an English-speaking academic audience, we excluded 3 articles not written in English. This ensures that the included studies are accessible and can be consistently interpreted without translation bias.
  4. 4.
    Filtering by Publication Type: To prioritize peer-reviewed journal articles, we excluded papers classified as Q3 or Q4 and proceedings papers, book chapters, editorials, and letters, removing 244 articles. This criterion ensures that our review is based on high-impact research published in top-quality journals, as lower-ranked journals or other types of publications may not always meet rigorous peer-review standards.
  5. 5.
    Checking for Relevance (Title, Abstract, and Conclusion Screening): The remaining articles were reviewed based on their title, abstract, and conclusions to assess their relevance to the research questions of this study. Articles that did not focus specifically on AI applications in SCRA were excluded. This step eliminated an additional 383 articles, leaving a final sample of 51 articles for in-depth review.
Each criterion was carefully chosen to ensure that only studies of high relevance and quality were included, thus minimizing potential biases. Using dual databases (Google Scholar and WoS) ensured comprehensive coverage, while the sequential exclusion criteria systematically filtered out irrelevant or low-quality studies. To demonstrate further transparency, Table 1 outlines the inclusion and exclusion criteria and their justifications.
3.4 Bias Mitigation
To address concerns about potential bias, we followed a systematic approach that minimizes subjective selection at each stage:
  1. 1.
    By using two widely recognized databases, Google Scholar and WoS, we ensured broad initial coverage of the literature on AI in SCRA, reducing the likelihood of selection bias by capturing a diverse range of studies.
  2. 2.
    Each exclusion criterion was pre-defined and systematically applied, following a step-by-step approach that minimized subjective judgment. This ensures that all articles are assessed based on consistent standards.
  3. 3.
    The PRISMA framework was applied rigorously to eliminate irrelevant studies at multiple stages, further reducing the risk of bias by ensuring that only studies meeting high-quality and relevance standards were retained.
3.5 Final Selection
After applying all criteria, 51 articles were deemed relevant and of high quality for inclusion in this review. The selected articles cover various AI methodologies applied to different SC risk scenarios, providing a comprehensive foundation for our analysis. According to Figure 1, 434 articles were included in the bibliometric analysis to provide a broad understanding of the field, considering various quantitative metrics. The 51 shortlisted articles represent a more focused subset. We selected them based on additional qualitative criteria, such as relevance, methodology, and contribution to the research.
Figure 1: Prisma flow diagram for systematic literature review.
3.6 Data Extraction and Synthesis
The selected articles underwent detailed analysis, focusing on methodologies, AI/ML techniques employed, key findings, and implications for SCRA. Data extraction forms were filled in to gather pertinent information for synthesis and further analysis. To understand the research landscape comprehensively, we performed bibliometric and cocitation analysis of the 434 articles resulting from the filtering process. The bibliometric analysis helped us identify the field’s most influential authors, journals, and key research themes. Cocitation analysis allowed us to identify clusters of related articles and determine their interconnections. Our study has several limitations that should be acknowledged. First, the search was limited to articles published between 2015 and 2024, and it is possible that relevant articles published before or after this timeframe were not included.
4 Bibliometric Analysis
To comprehensively analyze the scholarly landscape of SCRA using AI, we employed a combination of tools, including MS Excel and VOSviewer. This bibliometric analysis encompasses several dimensions, offering a deeper understanding of the field by identifying trends, contributions, and intellectual structures.
4.1 Publication Trend Over Time
Figure 2: Annual publication trend (2015-2024) of the reviewed literature.
Figure 2 illustrates the annual publication trends from 2015 to 2024. A significant rise in publications is observed post-2020, likely influenced by the COVID-19 pandemic, highlighting the critical need for resilient SC systems. The peak in research output during 2021 and 2022 reflects the heightened importance of AI/ML-driven risk assessment in tackling complex SC challenges. However, the slight decline in subsequent years may indicate a maturing research field, transitioning from exploratory studies to more application-oriented implementations in industry.
4.2 Geographical Distribution of Research
Figure 3: Geographical distribution of published papers.
Figure 3 visualizes the geographical distribution of research contributions. China leads with 164 publications, followed by the United States (74) and Australia (32). This dominance underscores the strong focus of Chinese institutions on AI-driven SCRA, supported by extensive investments in AI research and SC optimization. The global spread of contributions emphasizes the universal relevance of SCRA, with increasing participation from emerging economies.
Table 2: Top 10 authors according to the published papers
{tblr}
column2 = c, hline1-2,12 = -, Authors & Published Paper Count
Gupta S 5
Kumar A 4
Modgil S 4
Bag S 3
Hu XJ 3
Wang GJ 3
Xie C 3
Yao G 3
Zhu Y 3
Antoniou G 2
4.3 Analysis of Authors
As shown in Table  2, the top 10 authors in the field have significantly influenced the domain of SCRA using AI/ML. “Gupta S” leads with 10 publications, followed by “Kumar A” (8) and “Liu YK” (7). These authors represent thought leaders who have contributed extensively to developing AI-based risk assessment frameworks and methodologies.
4.4 Publisher Contributions
Figure 4: Publisher-wise distribution of the reviewed articles.
Figure 4 presents the contributions of various publishers. Elsevier is the most prominent, contributing over 120 articles, followed by Springer Nature and MDPI. These publishers serve as key platforms for disseminating high-quality research, reflecting their alignment with the evolving needs of AI and SCRA scholarship.
4.5 Journal Contributions
Figure 5: Journal-wise distribution of the reviewed articles.
In Figure 5, journal contributions are analyzed, with “Sustainability” being the most active source of publications, followed by “International Journal of Production Research” and “Annals of Operations Research.” These journals provide vital platforms for researchers to share AI-driven SC risk management advancements.
(a) Author co-citations network graph.
(b) Co-citation network graph of cited references.
(c) Co-citation network graph of publication sources.
Figure 6: Co-citation network graphs visualization using (a) authors, (b) cited references, and (c) sources as the unit of analysis.
4.6 Co-citation Analysis
Co-citation analysis was conducted to explore the intellectual structure of AI-SCRA research, focusing on authors, references, and publication sources. Figure 6(a) illustrates a network of 55 authors grouped into three clusters, emphasizing the pivotal roles of researchers such as Ivanov and Baryannis, who have significantly shaped the development of AI-driven risk assessment frameworks. Figure 6(b) focuses on cited references, identifying five distinct clusters highlighting recurring themes, including adopting hybrid AI models, developing advanced risk prediction frameworks, and exploring post-COVID SCR strategies. Lastly, Figure 6(c) underscores the centrality of academic journals, with the International Journal of Production Research standing out as a key publication venue, reflecting its substantial impact on the field’s scholarly discourse and research progression.
(a) Co-authorship network analysis of 48 authors.
(b) Co-authorship network analysis of 18 countries.
Figure 7: Co-authorship network graphs visualizing collaborations at (a) author and (b) country levels.
4.7 Co-authorship Analysis
Co-authorship analysis is a powerful method for understanding collaborations among authors and researchers based on joint publication activities. This analysis highlights the collaborative nature of the research domain, uncovering influential partnerships and the significant contributions of specific countries.
Figure 7(a) illustrates the co-authorship network among authors in this field. Using a threshold of at least one document and a minimum of 55 citations per author, 48 authors out of a total of 304 were identified as significant contributors. The visualization, generated using VOSViewer, reveals four distinct clusters, color-coded as red, green, blue, and yellow. These clusters represent groups of authors with strong collaborative ties. For example, the robust connections surrounding “Modgil” in the red cluster indicate frequent co-authorship and shared research initiatives, reflecting a high degree of collaboration within the cluster.
Figure 7(b) showcases the co-authorship network at the country level. By applying thresholds of at least two documents and nine citations per country, 18 out of 36 countries were included in the analysis. The visualization highlights England and China as leading contributors, dominating in terms of document count, citation impact, and total link strength. This dominance underscores their pivotal roles in advancing AI/ML-driven SCRA research. The results also emphasize strong international collaborations between these countries and other key players, fostering a globally interconnected research ecosystem.
(a) Co-occurrence network analysis of top 26 authors’ keywords.
(b) Co-occurrence network analysis of top 47 keywords used in the literature related to SCRA-AI.
Figure 8: The Co-occurrence Network visualizes the intricate relationships among identified keywords. Each node represents a keyword, and the edges denote the frequency of co-occurrence. Clusters of tightly connected nodes signify thematic groupings, providing a comprehensive snapshot of the intellectual structure within the analyzed literature.
4.8 Co-occurrence Analysis
The co-occurrence analysis offers valuable insights into the thematic structure of SCRA and AI/ML research by identifying how keywords are interconnected within the scholarly literature. Using VOSviewer, we conducted this analysis by focusing on the units of analysis, such as “Authors’ Keywords” and “Keywords Plus.” This approach reveals latent connections and thematic clusters, enabling a deeper understanding of this domain’s critical topics and prevalent themes.
The network map shown in Figure 8(a) illustrates the relationships among the 26 most frequently occurring author keywords. These keywords were selected from an initial pool of 1,483, applying a minimum threshold of three occurrences. Larger nodes, such as “artificial intelligence” and “machine learning,” indicate higher frequencies. The proximity of the nodes reflects their degree of association, with closely grouped nodes forming thematic clusters. Four clusters emerged, highlighting key themes, including risk management, sustainability, and digital transformation.
Figure 8(b) illustrates the co-occurrence of keywords found in the reviewed literature. The analysis identified 47 frequently occurring terms from a total of 1,032, using a minimum threshold of three occurrences. These keywords were categorized into five distinct clusters, reflecting prevalent themes and connections within the research field. Terms such as “management” and “framework” exhibited the strongest linkages, indicating their crucial roles in connecting different thematic areas. Additionally, other significant clusters highlighted emerging trends, including “supply chain resilience,” “blockchain,” and “sustainability.” This co-occurrence analysis offers a comprehensive overview of the intellectual structure of SCRA research, showcasing both established and emerging themes. The findings emphasize the interconnectivity of concepts and provide guidance for future research directions by identifying critical areas of scholarly focus.
Table 3: Top 10 Affiliations According to Paper Count
{tblr}
column2 = c, column3 = c, hline1-2,13 = -, Affiliations & Published Paper Count % of 51 articles
NEOMA BUSINESS SCHOOL 5 5.319
CHINESE ACADEMY OF SCIENCES 4 4.255
EMLYON BUSINESS SCHOOL 4 4.255
HEFEI UNIVERSITY OF TECHNOLOGY 4 4.255
HONG KONG POLYTECHNIC UNIVERSITY 4 4.255
HUNAN UNIVERSITY 4 4.255
INTERNATIONAL MANAGEMENT INSTITUTE (IMI) KOLKATA 4 4.255
UNIVERSITY OF CAMBRIDGE 4 4.255
UNIVERSITY OF TECHNOLOGY SYDNEY 4 4.255
UNIVERSITY OF HUDDERSFIELD 3 3.191
OTHERS 214 57.45
4.9 Top 10 Affiliations
Table  3 highlights the top 10 affiliations contributing to SCRA research, ranked by their published paper count. These institutions collectively demonstrate significant academic leadership, with NEOMA Business School leading with 5 publications (5.319% of the total reviewed articles). Several institutions, including the Chinese Academy of Sciences, Emlyon Business School, and Hong Kong Polytechnic University, each contributed 4 papers representing 4.255% of the reviewed literature. These affiliations underscore a global research effort, with prominent contributions spanning Europe, Asia, and Australia. Together, the top affiliations account for a substantial share of foundational studies in SCRA, promoting advancements in the application of AI to mitigate SC risks. Meanwhile, the category labeled “Others” collectively represents the remaining 57.45% of contributions, reflecting a broad and diverse research landscape.
4.10 Bibliographic Coupling
Bibliographic coupling assesses the similarity of documents, authors, or organizations based on shared references, revealing thematic alignments and intellectual structures. Key insights are as follows:
4.10.1 Authors
Figure 9 illustrates the bibliographic coupling analysis of 28 authors, grouping them into six thematic clusters based on shared references, highlighting intellectual connections in SCR research. Prominent contributors like Ivanov D, Gupta S, and Wang G-J demonstrate centrality, reflecting their significant influence in areas such as risk prediction, disruption management, and quantitative modeling. The clusters reveal interdisciplinary themes, including resilience strategies, industry-specific applications, and the integration of AI-driven methods. This analysis showcases both established research directions and emerging trends, providing a roadmap for future studies and collaborations in SCR.
Figure 9: Bibliographic coupling analysis of 28 authors.
4.10.2 Countries
Figure 10 illustrates the bibliographic coupling among 18 countries, categorizing them into clusters highlighting global research collaborations in AI-driven SCRA. Countries such as China, the United States, and Germany emerge as central hubs, showcasing their pivotal roles in advancing research and promoting international partnerships. The clusters signify thematic alignments across regions, with nations collaborating on critical topics like predictive analytics, sustainability, and resilience strategies. This visualization underscores the significance of cross-border collaboration in addressing global SC challenges through AI innovation.
Figure 10: Bibliographic coupling analysis of 18 countries.
4.10.3 Documents
Figure 11 illustrates the bibliographic coupling analysis of 62 documents, visually depicting the thematic clustering and citation relationships among the studies. The eight clusters highlight distinct research themes within the field, showcasing the interconnectedness of key works. Highly influential documents, such as those by “Baryannis G.” and “Ivanov D.,” emerge as central nodes, reflecting their significant contributions to the discourse on SCRA using AI. This visualization underscores the collaborative and evolving nature of the research domain, providing insights into how various studies align and contribute to specific thematic advancements.
Figure 11: Bibliographic coupling analysis of 62 documents.
4.10.4 Organizations
Figure 12 presents the bibliographic coupling analysis of 24 organizations, highlighting their collaborative and thematic alignments in AI-based SCRA research. The three clusters represent distinct institutional networks, with prominent contributors like the Hong Kong Polytechnic University and the Indian Institute of Technology Delhi playing central roles. This visualization emphasizes the global distribution of research efforts and showcases the leadership of key institutions in advancing AI-driven methodologies for SCRA.
Figure 12: Bibliographic coupling analysis of 24 organizations.
4.10.5 Journals
Figure 13 illustrates the bibliographic coupling analysis of the core academic journals that are driving the advancements in SCRA research by showing the interconnections and thematic clustering among 17 influential journals. Journals such as the “International Journal of Production Research” and “Annals of Operations Research” play a crucial role, providing foundational studies and methodologies that inform AI-driven approaches to SCRA. These clusters reveal thematic trends and emerging focus areas and facilitate knowledge dissemination by connecting researchers across disciplines.
Figure 13: Bibliographic coupling analysis of 17 publications sources.
Table 4: Publication Trends by Journal (2015-2024)
{tblr}
hline1-2,22 = -, Source Title & 2016 2017 2018 2019 2020 2021 2022 2023 Total
ANNALS OF OPERATIONS RESEARCH           2 4 1 7
COMPUTERS INDUSTRIAL ENGINEERING       1   1 2 2 6
INTERNATIONAL JOURNAL OF PRODUCTION RESEARCH       1     4 1 6
SUSTAINABILITY       1 1 1 1 1 5
WIRELESS COMMUNICATIONS MOBILE COMPUTING           1 3   4
EXPERT SYSTEMS WITH APPLICATIONS 1           1 1 3
INDUSTRIAL MANAGEMENT DATA SYSTEMS     1     1 1   3
INTERNATIONAL JOURNAL OF LOGISTICS MANAGEMENT           1 2   3
INTERNATIONAL JOURNAL OF PRODUCTION ECONOMICS       1 1     1 3
NEURAL COMPUTING APPLICATIONS   1       1 1   3
COMPUTATIONAL INTELLIGENCE AND NEUROSCIENCE             2   2
ENTERPRISE INFORMATION SYSTEMS         2       2
FUTURE GENERATION COMPUTER SYSTEMS       1     1   2
INTERNATIONAL JOURNAL OF DISTRIBUTED SENSOR NETWORKS       1 1       2
INTERNATIONAL JOURNAL OF OPERATIONS PRODUCTION MANAGEMENT             1 1 2
JOURNAL OF INTELLIGENT FUZZY SYSTEMS           2     2
KYBERNETES               2 2
OPERATIONS MANAGEMENT RESEARCH             2   2
ALEXANDRIA ENGINEERING JOURNAL             1   1
APPLIED SCIENCES-BASEL           1     1
4.11 Journal Publication Trends
Table 4 presents a comprehensive breakdown of published papers in various source titles from 2015 to 2024. These source titles encompass a range of journals and publications contributing to the field of SCRA. The table clearly shows the publication trends, indicating how many papers were published in each source title each year and their cumulative total. It provides valuable insights into the distribution of research output and the source titles that have been active in this area.
5 Framework for SCRA using AI
Figure 14: Framework for SCRA using AI.
Our proposed framework for AI-driven SCRA provides a comprehensive roadmap for SC professionals to integrate AI techniques effectively. Figure 14 visually shows the sequential steps involved, starting from defining key risk indicators to implementing preventive measures. By following these steps, practitioners can systematically define risk indicators, source and pre-process data, select and train models, and interpret results to facilitate informed decision-making. The framework is enriched with insights from existing studies on AI and ML applications in SC risk, presenting best practices, key considerations, and real-world challenges and opportunities identified in the literature.
5.1 Defining Key Risk Indicators
Our first step is identifying the key input and output variables that drive SCRA. Input variables include transactional data, supplier metrics, and external factors like political or environmental conditions, while output variables classify risks into severity levels such as “low,” “medium,” and “high” to align with organizational decision-making priorities. Defining these variables is crucial, as they represent the foundation upon which AI models assess risk. Wu, (2024) emphasize capturing operational and risk-related data for comprehensive risk assessments, with Xu et al., (2024) using historical sales and market trends for disruption prediction. Including financial and non-financial indicators, as highlighted by Wang et al., (2021), enhances the assessment of risk exposure. Practitioners should ensure these variables align with their industry’s risk characteristics to produce actionable and context-relevant insights.
5.2 Data Collection
Effective risk assessment relies on both internal and external data sources to capture a comprehensive picture of potential risks. Internal data from ERP systems, procurement records, and inventory management systems provides insights into operational processes. External data, such as market indicators, social media sentiment, and climate reports, are critical for Deiva Ganesh and Kalpana, (2022) used real-time data from social media, that is, Twitter, as the input source. ERP systems, supplier systems, and logistics systems, alongside external sources like social media and weather forecasts. Such data covers both structured (e.g., numerical data) and unstructured formats (e.g., text or multimedia) to enhance risk detection. Deiva Ganesh and Kalpana, (2022) used real-time data from social media, that is, Twitter, as the input source. The author used Twitter streaming API as a primary source for data extraction. Kosasih et al., (2022) utilized external platforms like Marklines for automotive SC data and Achilles Information Ltd for energy sector insights. Pan and Miao, (2023) collected data through surveys from industry practitioners to capture qualitative and contextual insights about SCR.Including such data enhances the model’s ability to preemptively detect risks like demand fluctuations or reputational issues, underscoring the value of a multimodal data approach.
5.3 Data Preprocessing
Preprocessing is a vital step. Raw data must undergo preprocessing to enhance model accuracy and efficiency. This step involves cleaning data to remove noise, normalizing values, and encoding categorical variables. Dimensionality reduction techniques like principal component analysis (PCA) applied by Luo et al., (2022) and imputation methods like residual mean replacement applied by Zhang et al., (2021) address challenges such as high-dimensional datasets and missing values. Time-series data, common in SC applications, requires additional adjustments like seasonality corrections. Techniques like Isolation Forest and Local Outlier Factor (LOF) detect anomalies, enhancing prediction reliability. SPSS statistical preprocessing methods have also been employed effectively in by Dang et al., (2022).
5.4 Model Selection
Practitioners have a range of model choices, each suited to different data characteristics and objectives in SCRA. Classical models like Support Vector Machine (SVM), Decision Tree (DT), and Logistic Regression(LR) are ideal for well-structured, low-dimensional datasets, where the data relationships are more straightforward and computational efficiency is prioritized. DL models, such as Convolutional Neural Networks (CNNs) and Long Short-Term Memory (LSTM) networks, are valuable for handling complex, high-dimensional data, including time-series demand data, where capturing temporal dependencies is essential. Hybrid models, which combine conventional machine learning (CML) and DL approaches, balance interpretability and accuracy, catering to scenarios where insights into model decisions are as important as predictive performance. Ensemble methods like Random Forest(RF) and XGBoost aggregate outputs from multiple models to enhance predictive accuracy and robustness. Additional research emphasizes that model selection should consider factors such as data availability, computational resources, and the necessity for model interpretability. Highly interpretable models are particularly important in regulated industries, such as finance, where transparency is critical. Hybrid and ensemble models, increasingly popular in recent literature, often effectively compromise predictive power and interpretability, making them suitable for complex SC risk scenarios.
5.5 Model Training and Hyperparameter Tuning
Model training typically involves cross-validation to improve model robustness. Tuning parameters like learning rate, depth, and number of estimators (in ensemble methods) is crucial to optimize performance. Tailored hyperparameter tuning can significantly enhance model performance in SC applications. Many researchers highlight the importance of automated tuning techniques, such as grid search, random search, and Bayesian optimization, to streamline model refinement and ensure that the final model is well-suited for specific risk scenarios. Wu et al., 2022b emphasized iterative optimization, while Zhang et al., (2022) addressed data imbalance with SMOTENC. Practitioners may also consider transfer learning approaches to leverage knowledge from similar datasets, especially in situations with limited labeled data.
5.6 Prediction
Once trained, the models generate predictions in the form of classifications (e.g., “high risk” vs. “low risk”), severity scores, or probabilities, which guide risk mitigation strategies. Baryannis et al., 2019b highlighted the role of predictive analytics in shielding SCs from disruptions. These predictions should ideally integrate into real-time dashboards to allow continuous monitoring and provide quick insights into risk fluctuations. This step facilitates proactive risk management, enabling organizations to adjust strategies promptly based on predicted risk levels and preventing disruptions before they escalate.
5.7 Interpretation of Results
Interpretability makes AI-driven predictions actionable, especially for stakeholders without a technical background. XAI, such as SHAP values or LIME, helps practitioners understand which variables most influence predictions. According to Olan et al., (2024), techniques like SHAP and LIME enhance decision-making by elucidating the rationale behind ML predictions, enabling better interpretability and fostering trust in SCRA. Interpretation tools also validate results with domain experts, ensuring the model’s output aligns with real-world SC conditions and providing a basis for actionable policy recommendations.
5.8 Preventive Measure
Based on the risk predictions and insights, practitioners can establish preventive strategies to mitigate potential SC disruptions. Common actions include adjusting inventory buffers, diversifying suppliers, or modifying production schedules. Preventive measures should be continuously refined based on AI predictions and risk evaluations. Incorporating feedback loops that allow practitioners to assess the impact of preventive measures on risk reduction is a good practice. Such loops can help organizations refine their approach over time, ensuring that the AI-driven risk assessment framework remains adaptable to changing conditions.
Table 5: Methods used and the key findings from the review articles.
Reference(s) Methods/Models Key Findings
Katsaliaki et al., (2022) Tensorization-based neural symbolic model Outperformed SNLP and GNN-LP in predicting relationships (AUC: 0.82, AP: 0.72).
Wang et al., (2022) Cost-sensitive learning to a RF (CSL-RF) model Improved prediction accuracy (Accuracy: 97.22%, AUC: 0.9850).
Wang et al., (2021) Adaptive heterogeneous multiview graph learning Improved forecasting accuracy (Accuracy: 87.44%).
Jianying et al., (2021) Optimized Backpropagation Neural Network (BPNN) and Particle Swarm Optimization-Backpropagation (PSO-BP) Outperformed the single Backpropagation (BP) (MAE: 2.1022, $R^{2}$: 0.932).
Chen et al., (2021) LSTM for forecasting LSTM model suitable for risk prediction (MAE: 1.037 - 1.52).
Zhu et al., (2016) RS-RAB Best prediction performance (Accuracy: 86.74%).
Bassiouni et al., (2023) Various DL approaches DL models improved accuracy by 14.46% (Softmax: 76.32%, SVM: 85.23%).
Zhu et al., (2019) RS-MultiBoosting superior performance of the RS-MultiBoosting ensemble method in forecasting
Athaudage et al., (2022) RF regression RF suitable for price prediction (MAE: 1.077, RMSE: 1.56, MAPE: 4.03%).
Liu and Huang, (2020) AdaBoosted SVM (EN-AdaPSVM) Outperforms other SVM models in noise data (Total False Rate: 11.76% (EN-AdaPSVM), 13.41% (AdaPSVM), 32.94% (CSVM)).
Zhu et al., (2017) Ensemble methods (RS-boosting and multi-boosting) RS-boosting outperforms other methods (Average Accuracy: 85.41% (RS-boosting), 84.08% (multi-boosting), 74.80% (boosting)).
Fu et al., (2022) BPNN, SVM, GA BP-GA model outperforms BP and SVM (Accuracy: BP-GA 97.19%, BP 82.33%, SVM 95.31%).
Zhang, (2022) Multimodal DL (DeepRisk) DeepRisk outperforms baseline methods.
Yang et al., (2021) Lasso-logistic model Lasso-logistic model has best prediction accuracy (Prediction Accuracy: 96.5%, Type II Error: 0.037).
Zhang et al., (2022) Fuzzy SVM (FSVM) FSVM outperform traditional SVM and DT (Higher accuracy in training and test sets).
Wang, (2021) Blockchain and fuzzy neural networks Average Accuracy: 84.11%, 84.66%(Independent Variables Set Vs).
Li, (2022) BPNN, LR BPNN and LR provide accurate predictions (Prediction Accuracy: 92.8%).
Bodendorf et al., (2023) DL model with causal inference High predictive performance and causal analysis.
Liu et al., (2022) Hybrid model (XGBoost–SMOTENC–RF) Development of a hybrid model chain, incorporating XGBoost, SMOTENC, and RF models (overall accuracy rate 91.67% and the average percentage error is 6.39%).
Rao and Li, (2022) LR, DT, integrated Logistic-RS Integrated RS improves prediction performance (Accuracy of Logistic: 0.831, Accuracy of Logistic-RS: 0.871, Accuracy of DT: 0.772, Accuracy of DT-RS: 0.795).
Handfield et al., (2020) Newsfeed analysis Predicted factory risks for a five-year planning horizon.
Pan and Miao, (2023) DL BPNN BPNN has high accuracy and low relative error (The maximum relative error calculated by the AHP is 57.41%, and the minimum is 5.88%, while the maximum relative error obtained by the BPNN is only 0.00210526%).
Wong et al., (2022) PLS-SEM, ANN AI-based models (AIRM) are significant determinants of RP and SCA.
Wang and Song, (2022) Interval Type 2 Fuzzy Neural Network (IT2FNN) IT2FNN punishes periodic deception, high accuracy (Accuracy 89%).
Nezamoddini et al., (2020) GA with NN Improved profits, lower inventory levels.
Xia et al., (2023) SVM, RF, MLP, LR RF showed best accuracy and low type I error.
Zhao and Li, (2022) SVM and BPNN Effective risk measurement, low relative error.
Li and Fu, (2022) PCA-GA-SVM PCA-GA-SVM outperforms SVM and GA-SVM, high accuracy.
Duan et al., (2021) BPNN optimized by GA Better fitting effect, higher prediction precision, and higher convergence speed.
Zhang et al., (2021) SVM optimized by firefly algorithm (FA-SVM) Improved classification efficiency and reduced error rates.
Luo et al., (2022) SVM, Particle Swarm Optimization, AdaBoost Improved classification effect, highest accuracy of 96.13%.
Dang et al., (2022) DL technology (BPNN) and blockchain Effective application of DL and blockchain.
Hosseini and Barker, (2016) Bayesian network (BN) BN allows disruption and improvement analysis.
Gao et al., (2022) Autoregressive integrated (AR), Mixture Density Networks (MDN) model, Autoregressive Integrated Moving Average (ARIMA) model, Multilayer Perceptron-Long short term memory (MLP-LSTM) model, Particle Swarm Optimization (PSO) algorithm and the improved PSO (IPSO) algorithm AR-MDN outperforms other models, IPSO has strong performance.
Wei, (2022) ML-based linear regression algorithm (ML-LRA) ML-LRA reduces supplier credit risk.
Zheng et al., (2023) Federated learning with dimensionality reduction (PCA) Collective risk prediction benefits organizations with small datasets (Maximum Difference: 0.0670 (FL vs. Local Learning), 0.0634 (FL vs. Centralized Learning).
Deiva Ganesh and Kalpana, (2022) Text-mining Use of social media data for contemporary risk understanding.
Zhang et al., (2023) Time-decayed long short-term memory (TD-LSTM), Deep Neural Network Improved predictability through TD-LSTM interpolation.
Baryannis et al., 2019a SVM and tree learning Framework offer good performance with interpretable and black-box methods (SVM Accuracy: 0.943, DT Accuracy: 0.950, Restricted DT (RDT) Accuracy: 0.916).
Han and Zhang, (2021) BPNN model BPNN effectively predicted high to extremely high risks in a food production company’s SC.
Belhadi et al., 2021b Hybrid EML approach: Gama Test (GT), Rotation Forest algorithm (RotF), Loogit Boosting algorithm (LB) Good performance of RotF-LB model (Best results for RotF-LB: Accuracy 91.72, Precision 92.96, Recall 95.53, F-measure 94.23, AUC 92.89).
Wu et al., 2022a BPNN and GA-BPNN Credit Risk Model Comparing iterations and risk prediction accuracy (GA-BPNN risk prediction accuracy above 0.92).
Lei et al., (2023) Binary image difference common area (BCAoID), SVM-SMA Highlight the effectiveness of the CGOA-SVM-SMA algorithm (Precision: 85.38%, F-score: 63%, TNR: 72%).
Yin et al., (2022) Convolutional neural network model (CNN) Optimal accuracy at 200 iterations (Comprehensive accuracy: 94.7%).
Janjua et al., (2023) Bi-LSTM CRF model Comparing Bi-LSTM CRF with Baseline CRF (the Bi-LSTM CRF model achieved an overall F1-Score of 85%, the Baseline CRF model achieved an overall F1-Score of 80%).
Yao et al., (2022) Hybrid model combining FS-MRI, SVME-AIR, and SMV AUC and KS scores for different parameters (AUC score of 0.8772 when the parameter “num_AIR” was set to 10, KS score of 0.6363 when the parameter “num_AIR” was set to 4).
Zhang et al., (2015) SVM and BPNN SVM model is superior to the BPNN model (BPNN Accuracy 66.67%, SVM model Accuracy 76.67%).
Aboutorab et al., (2022) RL-PRI approach Time efficiency of RL-PRI vs. manual process (Tarif Dispute Accuracy: 0.941463, Recession Accuracy: 0.907937, Sea Level Rise Accuracy: 0.964286, Flood Accuracy: 0.856322).
Wu, (2024) NLP techniques, word embedding models, bag-of-words approach Developed SCRisk measure, significantly correlated with operational buffers and stock return volatility.
Xu et al., (2024) Large language models (e.g., GPT-4, BERT) Improved decision support, optimized inventory management, reduced transportation costs, risk prediction.
Chen et al., (2024) GA-BPNN, Logistic Regression High accuracy in predicting MSME default and credit ratings (Accuracy: 92% for default prediction, 86% for credit rating).
6 Content Analysis of AI Techniques in SCRA
In recent years, the application of ML methods in SCRA has attracted substantial attention due to its potential to enhance decision-making processes and mitigate operational risks. This review paper explores various ML techniques applied in SCRA, highlighting their methodologies, key findings, and contributions.
6.1 Used ML Models and Key Findings
Table LABEL:tab:ml_methods summarizes various research methods or models used in risk prediction and analysis, along with the key findings and associated metrics from different research papers. Each row in the table represents a research paper’s first author, the research method or model used, and the specific findings and metrics achieved. These methods and findings are crucial in understanding risk prediction in various domains.
6.2 AI Models Used in SCRA
The application of AI in SCRA has seen significant growth, driven by the need for robust, predictive models to manage risk in complex and dynamic supply networks. To effectively analyze the methodologies employed in the reviewed literature, we classified the AI models into five main groups: CML Models, DL Models, Hybrid Models, Ensemble Methods, and Advanced Ensemble Methods. Each category is defined based on its distinct methodological characteristics, level of complexity, application scope, and suitability for specific SC risk scenarios, as shown in Figure 15.
Figure 15: Classification of AI models in SCRA based on methodology, complexity, and common application scenarios.
A detailed overview of each AI model used in SCRA, including their key characteristics, advantages, disadvantages, application cases, and suitable scenarios, is given below:
6.2.1 Conventional ML (CML) Models
Support Vector Machine (SVM): SVM is a supervised ML algorithm effective in high-dimensional spaces. It finds a hyperplane to separate classes in the feature space, often using kernel functions for non-linear separation. SVM performs well with limited data, making it useful for SCs with restricted historical data. By maximizing the margin, SVMs are less prone to overfitting, especially in smaller datasets. However, SVM can be computationally expensive for large datasets, which can be limiting in SCs with extensive data. Its results are also difficult to interpret, especially with complex kernel functions, which can be challenging for explainable predictions in SC risk. Bassiouni et al., (2023) found that the SVM classifier achieved an accuracy of 84.56% in predicting SC risks. Fu et al., (2022) found that the SVM model achieved a classification accuracy of 95.31% in assessing credit risk in Internet-based SC financing, outperforming the BPNN but slightly underperforming the BP-GA model. SVM is ideal for high-stakes, high-dimensional risk scenarios, like financial risk assessment with limited historical data. The data type is structured, often with limited samples. Input variables include supplier attributes, transaction history, and financial indicators, with the target variable being binary risk classification (e.g., high-risk vs. low-risk suppliers).
Decision Tree (DT): DTs are a tree-based algorithm that splits data into branches based on feature values, resulting in a flowchart-like model that is easy to interpret. DTs offer a simple structure that is easy to interpret, making it valuable in SC risk scenarios that require transparency. They are also relatively fast to train and execute, which makes them suitable for real-time or near-real-time assessments. However, DT models are prone to overfitting, especially with noisy data, and may require pruning or ensemble methods to mitigate this. They also struggle with capturing complex, non-linear relationships compared to models like neural networks. Rao and Li, (2022) found that the DT model achieved an accuracy of 77.2% in predicting SC financing risks. However, while highly interpretable, Baryannis et al., 2019a found that DT learning achieved lower performance predicting SC risks than SVM. The DT model demonstrated a trade-off between interpretability and performance, with SVM providing superior predictive accuracy, particularly in binary classification tasks for late deliveries. DT models are adequate for straightforward, interpretable assessments of SC risk, especially with structured and smaller datasets. The data type is structured with categorical or binary risk indicators. Input variables include financial ratios, transaction amounts, and supplier location, with the target variable being risk level classification (e.g., high, medium, low).
Logistic Regression (LR): LR is a linear model that estimates probabilities for categorical outcomes, commonly used for binary classification in risk assessment. LR offers a straightforward, easy-to-interpret linear approach that is beneficial for transparent decision-making in SC risk. It also requires less data and computational power, making it suitable for smaller datasets. However, LR is unsuited for complex, non-linear relationships, which can reduce accuracy in intricate SC risk scenarios. It also assumes feature independence, which might not hold in interconnected SC networks. Li, (2022) found that the LR model achieved a high prediction accuracy of 92.8% in forecasting investment risks in SCM, demonstrating excellent applicability and potential for broader use in SCM risk prediction. While Yang et al., (2021) found that the Logistic Regression model achieved a prediction accuracy of 83.1% in evaluating the credit risk of SMEs in SC finance, the lasso logistic model significantly outperformed it with an accuracy of 96.5% and a lower Type II error rate. LR is ideal for binary risk scenarios and cases where model transparency is essential. The data type is structured, typically tabular data. Input variables include supplier payment history, financial indicators, and supply volume, with the target variable being binary risk classification (e.g., default vs. no default).
6.2.2 Deep Learning (DL) Models
DL models, like neural networks, use multiple layers to model complex, non-linear relationships, making them practical for complex SC risk predictions. DL models are ideal for intricate, non-linear relationships in SC data, such as demand forecasting under uncertain conditions. They generally achieve high accuracy in well-established scenarios, particularly with large datasets. However, the “black-box” nature of DL models can make them difficult to interpret, which may limit adoption in risk-sensitive contexts. They also require large datasets and computational power, hindering some SC applications. Pan and Miao, (2023) found that the BPNN model demonstrated excellent performance in SCRA, achieving a maximum relative error of only 0.0021%, significantly outperforming the AHP model, which had a maximum relative error of 57.41%. This indicates the BPNN’s superior ability to evaluate risk accurately and reliably in SC systems. DL is suitable for large-scale, complex SCRAs with high data availability and less need for interpretability. The data type is large, structured, or unstructured data. Input variables include time series data on SC metrics, inventory levels, and demand trends, with the target variable being the probability of disruption or a risk score.
6.2.3 Hybrid Models (CML + DL Models)
Hybrid models combine CML and DL techniques to leverage the strengths of each for more robust SCRA. Hybrid models often outperform single models by combining strengths, such as using DL for feature extraction and CM for interpretability. They are versatile across various scenarios by tailoring components to specific data characteristics and risk types. However, hybrid models can be challenging to implement and tune, requiring expertise in multiple algorithms. Additionally, combining models can increase computational costs, which may be challenging for real-time applications. Fu et al., (2022) found that the hybrid BPNN and GA model outperformed both the standalone BPNN and SVM models in predicting credit risk for internet-based SC financing. The BPNN-SVM hybrid achieved the highest classification accuracy of 97.19%, significantly improving prediction performance compared to the BPNN (82.33%) and SVM (95.31%) models. Hybrid models are well-suited for applications requiring both high accuracy and moderate interpretability, especially when handling complex, multi-source data. The data type is mixed, including structured and semi-structured data. Input variables include supplier reliability, lead times, transactional volumes, and customer demand, with the target variable being multi-class risk categorization (e.g., high, medium, low).
6.2.4 Ensemble Methods
Ensemble methods like Bagging, Boosting, and Stacking combine multiple models to improve accuracy and robustness in SC risk prediction. Ensemble methods reduce variance and bias by aggregating multiple models, which makes them suitable for high-stakes risk assessment. They are versatile, as they can work with various algorithms, allowing for tailored approaches depending on the data and risk scenario. However, ensemble methods can require substantial computational resources, particularly for Boosting and Stacking. While they boost accuracy, they can decrease interpretability, making it difficult to understand individual model contributions. Liu and Huang, (2020) utilized the EN-AdaPSVM model, integrating AdaBoost with SVM and enhanced by noise elimination and adaptive mutation particle swarm optimization, achieving high classification accuracy and robustness in SC finance risk assessment. Luo et al., (2022) highlighted that while their AdaBoost-DPSO-SVM model showed significant performance improvements, it faced challenges such as higher error rates in negative sample recognition, suggesting potential issues with bias and misclassification. Ensemble methods are ideal for high-accuracy, high-stakes SC risk scenarios, especially when model interpretability is not the primary concern. The data type is structured, tabular data, often large datasets. Input variables include financial history, supplier performance scores, and geographic risk indicators, with the target variable being binary or multi-class risk levels (e.g., high-risk, low-risk, moderate risk).
6.2.5 Advanced Ensemble Methods
XGBoost: XGBoost is a powerful gradient-boosting algorithm known for its speed and accuracy, especially in structured data applications. XGBoost often outperforms traditional algorithms in structured data tasks, providing robust risk predictions. It also handles missing data well, which is useful in SC datasets that may be incomplete. However, XGBoost models can be complex, reducing interpretability and making tuning more challenging. Liu et al., (2022) developed a credit risk assessment model for the construction industry’s online SC finance using a hybrid approach combining XGBoost, SMOTENC, and RF. The study aimed to address the gaps in existing credit assessment systems, which often rely on traditional models with low accuracy and poor interpretability. Using data from 71 upstream enterprises between 2018 and 2020, the model achieved an overall accuracy rate of 91.67% on test samples. XGBoost is effective for structured, high-dimensional datasets in SC risk, particularly where interpretability is secondary to accuracy. The data type is large, structured data, often tabular. Input variables include financial indicators, transactional history, and geographic data, with the target variable being risk classification or score.
Random Forest (RF): RF is an ensemble method based on DTs, known for reducing variance and improving predictive accuracy. RF retains interpretability compared to more complex models and effectively handles high-dimensional data while being less prone to overfitting. However, RF can become slow with large datasets, particularly when building large forests. High-depth trees can also lead to overfitting if not carefully tuned. Wang et al., (2022) utilized a cost-sensitive Random Forest (CSL-RF) model to predict SMEs’ credit risk in SC finance, incorporating 40 diverse predictors and employing an imbalance sampling strategy for high accuracy. Athaudage et al., (2022) applied RF regression to forecast crude oil prices during disease outbreaks by integrating disease data with oil market variables, demonstrating reliable performance under complex conditions. Both studies highlight the effectiveness of RF models in handling varied data sets for robust forecasting. RF is effective when interpretability is moderately important and high accuracy and robustness are needed. The data type is structured, often large-scale. Input variables include supplier attributes, financial ratios, and location-based risk factors, with the target variable being multi-class risk level (e.g., high, medium, low).
7 Practical and Managerial Implications
The findings of this article will play a significant role in managerial strategies designed to enhance SCRM. First, it stresses that ML and AI are important tools that will give rise to a new phase of SCRA. Managers are encouraged to try advanced ML models, such as RF, XGBoost, and blends of these. They stimulate considerable risk assessment precision. These models are likely to enhance the accuracy of risk evaluation enormously and, thus, the effectiveness of their risk management strategies.
In addition, the study highlights the importance of flexibility and adaptability in an era after COVID-19, suggesting that companies need new SC strategies. Managers must now place contingency plans at the forefront, focusing on those not readily broken by any future disruptions. It is only by doing routine work thoroughly that one can concentrate on something special without fear or disturbance from the outside world, using people who have already developed a variety of fresh paths, so avoiding crisis after crisis with all that effort having been wasted.
The findings also highlight the importance of ensemble methods – RF and XGBoost – and hybrid methodologies in risk assessment. It is possible that such models could improve the accuracy of risk prediction and may be of interest to managers striving to strengthen their SCRM systems. The importance of time series prediction models, such as ARIMA and LSTM, cannot be overstated. Companies are advised to create tools predicting potential risks based on historical performance to approach the task of SC disruption mitigation with much-needed foresight.
XAI and ML models are extremely important. Models must produce precise and human-readable outputs as they assist decision-makers in understanding the process of risk assessment and even its findings. Furthermore, the study indicates the significance of real-time data mining and using non-traditional sources like social media, e.g., Twitter, as critical components of full-scale risk identification. Managers should be careful and up-to-date in understanding real-time information.
Blockchain represents an effective instrument for increasing transparency and traceability in the SC, thereby making risk monitoring mechanisms more efficient. Organizations exploring blockchain solutions should determine the possibilities of addressing and using the risk to boost SCR. SMEs require principal strategies in risk evaluation in terms of CR since they have distinct characteristics and risks in terms of operations. Conversely, implementing strategies based on SMEs’ needs and constraints is required. Moreover, federated learning should be implemented when data privacy issues arise. This method addresses the risk but protects confidential information, permitting partners to conduct joint risk forecasting. Compliance with security and data privacy must be performed almost immediately.
Several hybrid risk assessment methodologies that combine various ML models and techniques bring the necessary comprehensive view of SCRA, offering managers the chance to look into such approaches to develop risk management strategies. Additionally, comparing benchmarks of different ML models is critical to understanding more information about the best ones for particular risk scenarios. The overall message of the study is outlining the need for continuous performance evaluations and improvements. Risk assessment strategies should be able to evolve alongside other circumstances. Other assumptions refer to integrating ML-based risk assessment into already existing SC workflows and decision-making patterns. It is necessary to determine how such technologies can be easily implemented. Employee training in data analysis, ML, and AI is critical. Managers should invest in skill development to empower their SC professionals to harness these technologies effectively. Collaboration and knowledge sharing within organizations and across industries are encouraged to collectively build robust risk management strategies. Such partnerships have the potential to enhance risk resilience significantly.
Lastly, understanding and adhering to data protection and privacy regulations is critical when deploying AI and ML for risk assessment. Managers should ensure that their risk management practices align with the legal requirements to avoid compliance issues. The study offers a comprehensive roadmap for organizations aiming to fortify their SCRM strategies in an ever-evolving landscape.
8 Challenges and Limitations
Integrating AI in SCRA brings forth several challenges and limitations that researchers and practitioners must address. These challenges encompass the following:
  1. 1.
    Data Availability and Quality: AI models in SCRA rely heavily on data. However, ensuring the availability and quality of data remains a persistent challenge. SC data often comes from various sources, which can be fragmented, incomplete, or outdated. Researchers must explore ways to consolidate, cleanse, and augment data to enhance the effectiveness of AI models.
  2. 2.
    Interpretability and Explainability: Many AI techniques, such as DL and neural networks, are often considered “black-box” models. This lack of interpretability and explainability is a significant limitation. In SCRA, stakeholders require insights into why a model makes a particular prediction or decision. Developing AI models that are not only accurate but also interpretable is crucial for their adoption in real-world SCM.
  3. 3.
    Dynamic and Evolving Risk Factors: The SC environment is dynamic and continuously evolving. New risk factors emerge, and existing ones change over time. AI models may struggle to adapt to these dynamic conditions. Researchers need to explore techniques for real-time risk assessment and adaptive models that can respond to changing risk profiles.
  4. 4.
    Integration with Existing Systems: Integrating AI solutions into existing SCM systems can be challenging. Organizations often rely on legacy systems, and ensuring compatibility and seamless integration can pose hurdles. Research in this area should focus on creating modular AI solutions that can be integrated into diverse systems with minimal disruption.
9 Future Research Directions
In light of the challenges and limitations, future research directions in AI-based SCRA are multi-faceted and have the potential to advance the state-of-the-art. The development of AI models that deliver accurate predictions and offer comprehensive transparency and interoperability should be emphasized. Techniques like Local Interpretable Model-Agnostic Explanations (LIME) and SHapley Additive exPlanations (SHAP) serve as pivotal approaches to enhance the interpretability of AI models. By employing these methods, the goal is to demystify the decision-making process of AI models, enabling stakeholders to comprehend the rationale behind specific predictions or decisions. This transparency enhances trust and facilitates more effective integration of AI-driven insights into SC risk management strategies.
Combining multiple AI techniques, such as ML and expert systems, can enhance the robustness of SCRA. Future research should explore hybrid AI models that leverage the strengths of different approaches to improve prediction accuracy and reliability. Addressing data quality challenges involves multifaceted strategies. Developing robust methods for data cleansing and augmentation and ensuring high-quality data remains a critical focus. Beyond that, initiatives should be aimed at democratizing access to standardized SC data repositories. Creating and maintaining such repositories enables stakeholders across the SC spectrum to access reliable and consistent data. Increased accessibility fosters better-informed decision-making processes and promotes using standardized data sets for improved AI model training and validation. Investigating AI models capable of real-time risk assessment will be pivotal. These models should adapt to emerging risks and provide timely alerts or recommendations for risk mitigation.
Research on creating integration frameworks for AI-based SCRA is essential. These frameworks should facilitate easy integration into existing SC management systems, ensuring a smooth transition to AI-driven risk assessment. Expanding the application of AI in risk assessment to various domains within the SC, including logistics, procurement, and production, offers opportunities for comprehensive risk management. Investigate AI models capable of adapting to pandemic-induced disruptions within SCs. Focus on understanding the unique dynamics and challenges posed by COVID-19, aiming to develop AI-driven strategies that enable resilience and agility in such crisis scenarios. As AI plays an increasingly significant role in decision-making, researchers should also address ethical and legal considerations. This includes issues related to data privacy, bias, and fairness in AI models. By addressing these challenges and pursuing these research directions, AI-based SCRA can make significant strides toward more effective and reliable risk management in an ever-evolving global SC landscape.
10 Conclusions
This extensive review delved into the SCRA domain and examined the complex challenges of integrating AI, especially ML. Our thorough analysis encompassed a substantial pool of 1,439 SCRA papers, ultimately narrowing our focus to a refined subset of 51 papers. This endeavor has yielded significant insights and key findings, making substantial contributions to the SCRA literature.
Our primary contributions can be summarized by addressing the four key research questions:
  • RQ-1:
    Our review thoroughly analyzes the existing state of research concerning the application of AI and ML techniques in SCRM. We have meticulously assessed a substantial body of literature to extract meaningful insights, specifically 51 selected articles.
  • RQ-2:
    Through our analysis, we have identified the AI and ML techniques that are most commonly utilized in the domain of SCRA. This clarifies the methodologies employed and underscores the trends shaping the field.
  • RQ-3:
    Our review highlights the key findings and trends prevalent in the literature, shedding light on the progress, developments, and major areas of focus within the realm of AI and ML in SCRM.
  • RQ-4:
    We have identified critical research gaps and proposed future directions to guide further explorations in this field, ensuring the continued evolution of SCRA methods.
Despite these contributions, our research has some limitations. Our data primarily originated from Google Scholar and WoS rather than Scopus, potentially introducing biases. Our focus on papers published between 2015 and 2024 presents a limitation, as it may omit earlier relevant research that could provide valuable historical context. Additionally, our restriction to English-language papers may have omitted valuable non-English research.
Looking forward, several promising avenues for future research emerge in the domain of SCRA using AI. These directions encompass operationalizing interconnectedness, transformability, and sharing within SCRA frameworks, investigating evolving ICT roles in prediction and response, revisiting the cost implications of resilience enhancement, and exploring emerging technologies like blockchain. These endeavors are expected to enhance risk assessment effectiveness in the post-COVID era. Reviewers should adopt a comprehensive approach to gain deeper insights into this field by expanding their search to diverse databases, including non-English and grey literature sources, using snowballing techniques, collaborating with experts from related disciplines, leveraging ML tools for data analysis, and staying up-to-date with the latest research. Combining various search methods and expert opinions, such a multidisciplinary strategy is essential for uncovering valuable insights and emerging trends in this dynamic and critical field.
Our SLR, driven by these four primary research questions, has contributed valuable insights into the AI/ML application in SCRA, identified common techniques, outlined key findings and trends, and proposed essential research directions. This work serves as a guide for both researchers and practitioners, facilitating advancements in SCRM.